Scott and Martha work in a software company. Scott can make 10 schedules an hour, or he can perform 30 quality assurance checks an hour. On the other hand, Martha can make 8 schedules an hour or perform 28 quality assurance checks an hour. Who has the comparative advantage for making schedules, Scott or Martha? When you post your response, make sure you answer these two questions: . Who has the comparative advantage? . What are each persons opportunity costs? To complete your answer, youll need to make up a new story involving two characters producing goods. Post it along with your answers to the questions.
Answer #1

Making schedules (in one hour)

Quality assurances checks (in one hour)

Scott

10

30

Martha

8

28

Opportunity cost of making schedules in terms of giving up quality assurance checks

Opportunity cost of making quality assurance checks in terms of giving up making schedules

Scott

30/10=3

10/30=0.33

Martha

28/8=3.5

8/28=0.27

For Scott the opportunity cost of making schedules in terms of giving up quality assurance checks is 3 which is lower than 3.5 for Martha.

For Martha the opportunity cost of making quality checks in terms of giving up making schedules is 2.7 which is lower than 0.33 for Scott.

Scott has a lower opportunity cost of making schedules in terms of giving up quality assurance checks so Scott has a comparative advantage in making schedules.

Martha has a lower opportunity cost of making quality assurance checks in terms of making schedules so Martha has a comparative advantage in making quality checks.